# Large positive magnetoconductance in carbon nanoscrolls

### v1

@@ Abstract
We theoretically demonstrate that carbon nanoscrolls – spirally wrapped graphene layers with open endpoints – can be characterized by a large positive magnetoconductance. We show that when a carbon nanoscroll is subject to an axial magnetic field of $\simeq 10$T, the ballistic conductance at low carrier densities of the nanoscroll has an increase of about 200%. Importantly, we find that this positive magnetoconductance is not only preserved but can be even enhanced in the presence of on-site disorder. We prove that the positive magnetoconductance comes about the emergence of magnetic field-induced zero energy modes, specific of rolled-up geometries. Our results establish curved graphene systems as a new material platform displaying sizable magnetoresistive phenomena.
@@ 1 Introduction
Magnetocondutance – the change of conductance in response to an externally applied magnetic field – appears in different magnetic and non-magnetic materials alike and can have various physical origins. At very low temperatures, the presence of quantum interference effects, specifically weak (anti)localization, leads to a positive (negative) magnetoconductivity 1. Weak antilocalization has been recently observed for instance in topological insulators 2 and is related to the strongly spin-orbit coupled Dirac surface states of these materials. In the two-dimensional electron gas formed at LaAlO3/SrTiO3 interfaces, a combination of spin-orbit coupling and scattering by finite-range impurities gives rise to a single particle mechanism of positive magnetoconductance in response to in-plane magnetic fields and at temperatures up to the 20K range 3. Furthermore, the negative longitudinal magnetoresistance of Weyl semimetals 4 is connected to the chiral anomaly of Weyl fermions. It can reach values up to $40\%$ and exhibits a strong angular dependence 5, 6.
In nanostructures, geometrical effects owing to atomic structures, 7, 8 strain- and defect-engineering 9 or layer stacking 10 can be the platform for magnetoresistive phenomena as well. In topological insulators nanowires, for instance, the $\pi$ Berry phase due to the spin-momentum locking of the surface states leaves its hallmark on the electronic band structure and provides a gap in the energy spectrum 11. When threaded by a half magnetic flux quantum, the surface state gap effectively vanishes 12, thereby implying a positive magnetoconductance 13, 14 and Aharonov-Bohm oscillations. Beyond the linear transport regime, geometrical effects have been also shown to lead to dipolar distributions of Berry curvature 15, 16, 17 and consequently to the observation of a non-linear Hall effect in the presence of time-reversal symmetry 15.
In this study, we focus on the magnetotransport properties of carbon nanoscrolls 18. This compact nanoarchitecture can be synthesized by rolled-up technology and can be seen as radial superlattices due to their spiral cross section. This results in a very peculiar bandstructure and transport behavior different from conventional flat nanostructures 19. Both blue phosphorous 20 and black phosphorus nanoscrolls 21 are characterized by high carrier mobility. Importantly, there has been growing attention on aluminum- and lithium-based batteries that make use of carbon-based radial-superlattice cathodes 22, 23, 24, 25, 26.
@@
Figure 1: The conductances of two-winding CNSs with and without applied magnetic fields. (a) The way of applying magnetic fields and the boundary conditions of interface. (b) The conductances for the case of without disorder. (c) The conductances for the case of with disorder. In (b) and (c), the red line indicates the applied magnetic field $B=0$ Tesla, whereas the blue line is for $B\approx 10$ Tesla (10.3949 Tesla in the numerical calculation).
@@ 2 Methods
In order to analyze ballistic transport in carbon nanoscrolls (CNS), we employ both a continuum ${\bf k\cdot p}$ model 28 and a tight-binding model, with which we perform numerical calculations using the Kwant package 29. In the following, we consider a two-winding CNS that can be mapped to bilayer graphene (see Fig. 1(a)) with mixed boundary conditions 28. The corresponding four-band continuum model takes into account the sublattice and layer degrees of freedom and can be written in the $A1,B1,A2,B2$ basis 30. The resulting energy dispersion can be obtained from the relation $\hbar vk_{\pm}=\sqrt{\varepsilon^{2}\pm\gamma_{1}\varepsilon-\hbar^{2}v^{2}k_{z}^{2}}$, where $k_{\pm}$ and $k_{z}$ are the momenta in the tangential and axial direction of the CNS respectively. In the equation above we introduce the velocity $v=\sqrt{3}a\gamma_{0}/2\hbar$ (see Section S2 in Supporting Information). Additionally, we fix the intralayer coupling strength between sites A1 (A2) and B1 (B2) at $\gamma_{0}=3.16$ eV, and the interlayer coupling strength between sites A2 and B1 at $\gamma_{1}=0.381$ eV. We also consider a Möbius tube geometry obtained by closing the open boundaries of the two-winding CNS.
We define a positive magnetoconductance coefficient (PMC) as
@@
where $G(\phi)$ indicates the conductance with magnetic flux $\phi$. The two-terminal conductance in the ballistic regime is given by the Landauer formula 31, 32, 33, 28,
@@
where $f$ is the Fermi-Dirac distribution function, and $E_{F}$ is the Fermi energy. The zero temperature perfectly ballistic conductance of our one-dimensional nanostructure is proportional to the number of modes and given by $G(E,0)=2e^{2}N_{s}/h$. We neglect the mild spin-orbit coupling of graphene.
To account for the effect of disorder, we include a random on-site potential that is Gaussian distributed 28, 34 (see Section S3 in Supporting Information). We consider two characteristic disorder strengths of $0.1$ eV and $0.5$ eV respectively, both comparable to the intralayer hopping amplitude. The conductance is averaged over 10 random disorder configurations.
@@ 3 Results and discussion
Figure 2: The conductances of ribbon and two-winding CNSs for different disorder strengths. From (a) to (c), the red and blue line indicate the applied disorder $0.1$ eV and $0.5$ eV, respectively. The gray dashed line denotes the conductance for the perfect lattice. (a) The conductances of ribbon with no interlayer coupling strength, i.e. $\gamma_{1}=0$ eV. (b) The conductances of two-winding CNSs without applied magnetic fields. (c) The conductances of two-winding CNSs with applied magnetic fields $B\approx 10$ Tesla (10.3949 Tesla in the numerical calculation). As the counterparts of (a) to (c), (d) to (f) shows the conductance with the x-axis resenting energies.
@@
To get a comprehensive understanding of the transport properties of a two-winding CNS threaded by a magnetic flux, we first study the two-terminal conductance of a monolayer graphene nanoribbon with zigzag edges and ribbon width equal to the total arclength of our two-winding nanoscroll. Fig.  Figs. 2(a) and (d) show that in the low carrier density regime (electron or hole carrier density lower than 0.015 nm-2) , and thus close to the charge neutrality point (Fermi energy $|E|<30$ meV), the ballistic conductance is simply given by $G_{0}$. Disorder leads to a zero-conductance dips close to the charge neutrality point (see red and blue lines in  Figs. 2(a) and (d)). Very similar features are encountered when considering a two-winding CNS in the absence of externally applied fields (see Figs. 2(b) and (e)). We note that as compared to their flat counterpart CNSs display an ideal ballistic conductance that can be tripled even in the low carrier density regime, due to a multiplication of the number of modes. Additionally, the zero-conductance dip is narrower thus suggesting that conduction in curved graphene nanostructures is more robust against disorder. For a CNS threaded by a half-integer magnetic flux quantum, the situation is completely different. As shown in  Figs. 2(c) and (f), the ballistic conductance at charge neutrality is tripled. Furthermore, adding disorder does not lead to any zero-conductance dip even for disorder strength of about $0.5$eV, and thus larger than the interlayer hopping amplitude (see the blue line in  Figs. 2(c) and (f)). We thus find that a CNS is characterized by a PMC that reaches 200% close to the charge neutrality point.
@@
The conductance tripling in CNS threaded by a half-integer magnetic flux quantum can be understood by considering the electronic characteristic of CNSs and we start by considering a Möbius-like geometry in which the open endpoints of the CNS are closed (see Fig. 3(a)). Close to the ${\bf K}$ ($K^{\prime}$) valley we observe the appearance of a two-fold degenerate zero-energy modes. This zero energy modes disappear when opening the boundary conditions as in an actual CNS (see  Figs. 3 (b)). Instead, we observe the appearance of the characteristic zero-energy edge modes of zigzag terminated graphene. With a half-integer magnetic flux quantum the energy spectrum for a Möbius-like CNS does not qualitatively change – we only observe a shift in the axial momentum of the doubly degenerate zero energy modes (see Fig. 3(c)). The case of a CNS with open boundary conditions threaded by a magnetic flux retains the zero-energy zigzag edge states found in the absence of magnetic fields. However, we concomitantly find the emergence of the zero energy doublet found with closed boundary conditions (see Fig. 3(d)). It is the appearance of these additional modes that lead to the tripling of the ballistic conductance in the vicinity of charge neutrality point (the region of low carrier densities). To identify the qualitative difference between the Möbius-like-CNS and CNS ground states, the charge density distributions of ground states in the CNS with and without magnetic flux are both provided in  Fig. S3 in Supporting Information.
@@
Figure 3: Energy bands of the two-winding CNSs and Möbius tube for the K point: (a) Möbius tube and (b) two-winding CNSs which are without applied magnetic fields, whereas (c) Möbius tube and (d) two-winding CNSs are with applied magnetic fields $B\approx 10$ Tesla (10.3949 Tesla in the numerical calculation)
@@ 4 Conclusion
To sum up, we have theoretically demonstrated that radial superlattices, especially in AB- and BA stacked domain wall featuring two-winding CNS with magnetic flux, display a giant PMC. We have found that the PMC of a two-winding CNS is giant and up to more than two times of that of the ordinary graphene nanoribbon. With simulations of disordered systems, we have found that its conductance is less prone to disorder and PMC even increases, in contrast to the disordered TI nanowire that PMC decreases remarkably 11, 12.
To interpret this novel result, we gave the model of the Möbius tube and compared the band structures with and without magnetic flux at the very low energy near $E=0$. We show that the geometry of the radial superlattices, including the boundary conditions, interlayer coupling, and magnetic flux are critical in electronic transport. Our results afford new and pivotal insights to create and artificially control the high-disorder resistant Möbius tube ground state and edge states in curved carbon nanomembrane. It is expected that the insights and effects we have unveiled in our work will be observed in the experimental field.

### v3

@@ Abstract
We theoretically demonstrate that carbon nanoscrolls – spirally wrapped graphene layers with open endpoints – can be characterized by a large positive magnetoconductance. We show that when a carbon nanoscroll is subject to an axial magnetic field of several Tesla, the ballistic conductance at low carrier densities of the nanoscroll has an increase of about 200%. Importantly, we find that this positive magnetoconductance is not only preserved in an imperfect nanoscroll (with disorder or mild inter-turn misalignment) but can even be enhanced in the presence of on-site disorder. We prove that the positive magnetoconductance comes about the emergence of magnetic field-induced zero energy modes, specific of rolled-up geometries. Our results establish curved graphene systems as a new material platform displaying sizable magnetoresistive phenomena.
keywords:
Radial Superlattice, Aharonov-Bohm Effect, Interfacial Magnetic States, Longitudinal Magnetic Field
@@ 1 Introduction
Magnetocondutance – the change of conductance in response to an externally applied magnetic field – appears in different magnetic and non-magnetic materials alike and can have various physical origins. At very low temperatures, the presence of quantum interference effects, specifically weak (anti)localization, leads to a positive (negative) magnetoconductivity 1. Weak antilocalization has been recently observed for instance in topological insulators 2 and is related to the strongly spin-orbit coupled Dirac surface states of these materials. The competition between weak localization and antilocalization in InGaAs-based two-dimensional (2D) systems was analyzed through magnetoconductance. 3 In the 2D electron gas formed at LaAlO3/SrTiO3 interfaces, a combination of spin-orbit coupling and scattering by finite-range impurities gives rise to a single particle mechanism of positive magnetoconductance in response to in-plane magnetic fields and at temperatures up to the 20K range 4. Furthermore, the negative longitudinal magnetoresistance of Weyl semimetals 5 is connected to the chiral anomaly of Weyl fermions. It can reach values up to $40\%$ and exhibits a strong angular dependence 6, 7.
In nanostructures, geometrical effects owing to atomic structures, 8, 9 strain- and defect-engineering 10 or layer stacking 11 can be the platform for magnetoresistive phenomena as well  12. For instance, the ballistic magnetoconductance calculated in carbon nanotube reveals a step-like structure as a function of magnetic flux. 13 In topological insulators (TI) nanowires, the $\pi$ Berry phase due to the spin-momentum locking of the surface states leaves its hallmark on the electronic band structure and provides a gap in the energy spectrum 14. When threaded by a half magnetic flux quantum, the surface state gap effectively vanishes 15, thereby implying a positive magnetoconductance 16, 17 and Aharonov-Bohm oscillations. Additionally, magnetotransport has been theoretically studied in shaped TI nanowires, such as nanocones and dumbbells 18, 19, where the surface electrons experience an out-of-plane component of the coaxial magnetic field. This variation in cross-sectional area leads to unconventional magnetic transport properties. Furthermore, geometrical effects have been also shown to lead to dipolar distributions of Berry curvature 20, 21, 22 and consequently to the observation of a non-linear Hall effect in the presence of time-reversal symmetry 20.
In this study, we focus on the magnetotransport properties of carbon nanoscrolls with a turn number of two or fewer  23, 24, 25, 26. This compact nanoarchitecture can be synthesized by rolled-up technology and can be seen as radial superlattices due to their spiral cross section. This results in a very peculiar bandstructure and transport behavior different from conventional flat nanostructures 27, 28. Both blue phosphorous 29 and black phosphorus nanoscrolls 30 are characterized by high carrier mobility. Importantly, there has been growing attention on aluminum- and lithium-based batteries that make use of carbon-based radial-superlattice cathodes 31, 32, 33, 34, 35.
@@
Figure 1: The conductances of two-turn CNSs with and without applied magnetic fields. (a) The way of applying magnetic fields and the boundary conditions of interface. (b) The conductances for the case of without disorder. (c) The conductances for the case of with disorder. In (b) and (c), the red line indicates the applied magnetic field $B=0$ Tesla, whereas the blue line is for $B\approx 10$ Tesla (10.3949 Tesla in the numerical calculation).
@@ 2 Methods
In order to analyze ballistic transport in carbon nanoscrolls (CNS), we employ both a continuum ${\bf k\cdot p}$ model 37 and a tight-binding model, with which we perform numerical calculations using the Kwant package 38. In the following, we consider a two-turn CNS that can be mapped to bilayer graphene (see Fig. 1(a)) with mixed boundary conditions 37. The corresponding four-band continuum model takes into account the sublattice and layer degrees of freedom and can be written in the $A1,B1,A2,B2$ basis 39. The resulting energy dispersion can be obtained from the relation $\hbar vk_{\pm}=\sqrt{\varepsilon^{2}\pm\gamma_{1}\varepsilon-\hbar^{2}v^{2}k_{z}^{2}}$, where $k_{\pm}$ and $k_{z}$ are the momenta in the tangential and axial direction of the CNS respectively. In the equation above we introduce the velocity $v=\sqrt{3}a\gamma_{0}/2\hbar$ with lattice constant $a$ (see Section S1 in Supporting Information).
We construct a tight-binding model for a carbon nanoscroll by rolling a zigzag graphene nanoribbon perpendicular to its edges, restricting to AB-stacked (Bernal stacked) structures. The model accounts for nearest-neighbor and interlayer hoppings. The unit cell consists of pairs of A-B carbon atoms from the nanoribbon, represented as $;\{A_{1},B_{1};A_{2},B_{2};$$\ldots;A_{m},B_{m}\}$, as shown in Fig. 2(a), where $m$ is the number of pairs of A-B carbon atoms in the unit cell. Along the zigzag boundaries, the carbon atoms of different layers are also aligned according to the AB-stacking configuration. To form a two-turn CNS with an AB-stacked structure, an example with 7 pairs of A-B carbon atoms ($m=7$) is presented in  Figs. 2(b) and (c) 40, 41. In this structure, half of the atoms sit above the centers of the hexagons, while the others are directly above the atoms of the inner layer.
Figure 2: (a) The unit cell, as denoted by green rectangular, of CNS is defined along the axis perpendicular to the core axis and has open boundaries. The length of the unit cell is determined by the arclength of the nanoscroll. (b) The cross section of a CNS featuring two turns. (c) The flattened interlayer structure of the CNS shows an AB-stacking arrangement. The dashed lines in (b) and (c) denote the interlayer coupling $\gamma_{1}$.
For modeling the two-turn CNS, we fix the intralayer coupling strength between A and B sites at $\gamma_{0}=3.16$ eV, and the interlayer coupling strength between site A2 (A site in the 2nd turn) and site B1 (B site in the 1st turn) with $\gamma_{1}=0.381$ eV, corresponding to an interlayer distance of $3.35$  in the AB-stacked bilayer graphene. 42, 39, 43 The lattice constant $a$ is 2.4595 Å 44 with carbon-carbon bond length 1.42 Å for graphene. For the system length scale, we set the total arc length to $X=100$ nm, which contains 934 atoms, for both the two-turn nanoscroll and the Möbius tube in the Kwant simulation. 45 This corresponds to a perimeter of $L=50$ nm and a radius of $7.99$ nm for a single turn. Furthermore, the length along the core axis is $300$ nm.
We define the positive magnetoconductance coefficient (PMC) as
@@
where $G(\phi)$ indicates the conductance with magnetic flux $\phi$. The two-terminal conductance in the ballistic regime is given by the Landauer formula 46, 47, 48, 37,
@@
where $f$ is the Fermi-Dirac distribution function, and $E_{F}$ is the Fermi energy. The zero temperature perfectly ballistic conductance of our one-dimensional nanostructure is proportional to the number of modes ($N_{s}$) and given by $G(E,0)=2e^{2}N_{s}/h$. We neglect the mild spin-orbit coupling of graphene.
To account for the effect of disorder, we include a random on-site potential that is Gaussian distributed 37, 49 (see Section S2 in Supporting Information). We consider two characteristic disorder strengths of $0.1$ eV and $0.5$ eV respectively, both comparable to the intralayer hopping amplitude. We examined the convergence of the averaged conductance and found that 200 configurations already achieve a small fluctuation of $5\%$. Therefore, we use 200 random disorder configurations for the results presented in the article, unless otherwise stated. More details on disorder convergence tests and, in addition, the calculations for the localization length, proportional to the mean free path in a (quasi-) one-dimensional system, 50 are provided in Section S2 of the Supporting Information.
@@ 3 Results and discussion
Figure 3: The conductances of (a) monolayer ribbon, (b) bilayer ribbon, (c) two-turn CNS and (d) two-turn CNS with applied magnetic flux for different disorder strengths. (e) The conductances of a monolayer ribbon, i.e. $\gamma_{1}=0$ eV. (f) The conductances of a AB-stacked bilayer ribbon with interlayer coupling strength $\gamma_{1}=0.381$ eV. (g) The conductances of a two-turn CNSs without applied magnetic fields. (h) The conductances of two-turn CNSs with applied magnetic fields $B\approx 10$ Tesla (10.3949 Tesla in the numerical calculation). As the counterparts of (e) to (h), (i) to (l) shows the conductance with the x-axis resenting energies. The gray dashed line denotes the conductance for the perfect lattice. The red and blue line indicates the result of a system with applied disorder $0.1$ eV and $0.5$ eV, respectively.
@@
To get a comprehensive understanding of the transport properties of a two-turn CNS threaded by a magnetic flux, we first study the two-terminal conductance of a monolayer graphene nanoribbon with zigzag edges and ribbon width equal to the total arclength of our two-turn nanoscroll (see Fig. 3(a)). Based on order-of-magnitude estimation, we expect that the Zeeman effect and spin-orbit coupling have a negligible impact on the large PMC 51, 52.  Figures 3(e) and (i) show that in the low carrier density regime (electron or hole carrier density lower than 0.015 nm-2) , and thus close to the charge neutrality point (Fermi energy $|E|<30$ meV), the ballistic conductance is simply given by $G_{0}$. Disorder leads to a zero-conductance dips close to the charge neutrality point (see red and blue lines in Figs. 3(e) and (i)). Very similar features are encountered when considering either a bilayer graphene ribbon (see Figs. 3(f) and (j)), a two-turn CNS in the absence of externally applied fields (see Figs. 3(g) and (k)), or a double-walled carbon nanotube (see Section S3 in Supporting Information).
For a CNS threaded by a half-integer magnetic flux quantum (see Fig. 3(d)), the situation is completely different. As shown in  Figs. 3(h) and (l), the ballistic conductance at charge neutrality is tripled, comparing to the results of monolayer nanoribbon shown in Figs. 3(e) and (i). Furthermore, adding disorder does not lead to any zero-conductance dip even for disorder strength of about $0.5$ eV, and thus larger than the interlayer hopping amplitude (see the blue line in Figs. 3 (h) and (l)). We thus find that a CNS is characterized by a PMC that reaches 200% near the charge neutrality point. Moreover, the localization length along the core axis of a two-turn CNS with a magnetic flux exceeds one micrometer, as detailed in Section S2 of the Supporting Information. This indicates that PMC can be realized in nanoscroll systems with length scales ranging from nanometers to micrometers 23, 26.
We note that the additional phase of the nanoscorll states, determined by the applied magnetic flux, is given by $\phi=\pi(\frac{L}{2\pi})^{2}B$, where $B$ is the magnetic field strength and $L$ is the one-turn length of the nanoscroll. For our proposed magnetotransport to occur at $\phi=\pi/2=\pi(L/2\pi)^{2}B_{c}$, the required magnetic field strength $B_{c}$ can be reduced by a factor of $N^{2}$ times by increasing the nanoscroll’s arclength by a factor of $N$. For a two-turn nanoscroll with an arclength eof $150$ nm, for example, the required field strength, achieving the results shown in Figs. 3(h) and (l), can be reduced to approximately 4.6 Tesla (see Section S3 in Supporting Information). Additionally, the energy and conductance under various applied magnetic fields are presented in Section S4 of the Supporting Information.
@@
The conductance tripling in CNS threaded by a half-integer magnetic flux quantum can be understood by considering the electronic characteristic of CNSs. We start by considering a Möbius-like geometry in which the open endpoints of the CNS are closed (see Fig. 4(a)). Close to the ${\bf K}$ ($K^{\prime}$) valley we observe the appearance of a two-fold degenerate zero-energy modes. This zero energy modes disappear when opening the boundary conditions as in an actual CNS (see Fig. 4(b)). Instead, we observe the appearance of the characteristic zero-energy edge modes of zigzag terminated graphene. With a half-integer magnetic flux quantum the energy spectrum for a Möbius-like CNS does not qualitatively change – we only observe a shift in the axial momentum of the doubly degenerate zero energy modes (see Fig. 4(c)). The case of a CNS with open boundary conditions threaded by a magnetic flux retains the zero-energy zigzag edge states found in the absence of magnetic fields. However, we concomitantly find the emergence of the zero energy doublet found with closed boundary conditions (see Fig. 4(d)). It is the appearance of these additional modes that lead to the tripling of the ballistic conductance in the vicinity of charge neutrality point (the region of low carrier densities).
To further demonstrate that the doubly degenerate states at zero energy in the CNS with a magnetic flux are inherited from the nontrivial interfacial states in the Möbius-like CNS, we have estimated the charge density distributions of the zero-energy states in the Möbius-like CNS, the Möbius-like CNS with an applied magnetic flux, and the CNS with the same magnetic flux. The results, shown in Fig. 5, confirm this connectionWe would like to emphasize that pioneering studies 53, 54, 55, 56 have shown that the AB-BA interface in bilayer graphene induces a topological feature in $k$-space, resulting in 1D interfacial topological valley states. Our findings in Figs. 4 and 5 demonstrate that the nontrivial interfacial state is sustained not only in a Möbius tube with the same interface but also in a carbon nanoscroll under an applied magnetic field.
@@
Figure 4: Energy bands of the two-turn CNSs and Möbius tube for the K point: (a) Möbius tube and (b) two-turn CNSs which are without applied magnetic fields, whereas (c) Möbius tube and (d) two-turn CNSs are with applied magnetic fields $B\approx 10$ Tesla (10.3949 Tesla in the numerical calculation)
Figure 5: The distributions of charge densities of the doubly degenerate zero energy states in: (a) Möbius tube, (b) Möbius tube with applied magnetic flux, and (c) CNS with applied magnetic field. The results are obtained by calculating tight-binding models of the Möbius tube and the CNS.
@@ 4 Conclusion
To sum up, we have theoretically demonstrated that radial superlattices, especially in AB- and BA stacked domain wall featuring two-turn CNS with magnetic flux, display a giant PMC. We have found that the PMC of a two-turn CNS is giant and up to more than two times of that of the ordinary graphene nanoribbon. With simulations of disordered systems, we have found that its conductance is less prone to disorder and PMC even increases, in contrast to the disordered TI nanowire that PMC decreases remarkably 14, 15.
To interpret this novel result, we developed a model of the Möbius tube with an AB-BA bilayer interface and compared its band structures and quantum states with and without magnetic flux. The proposed PMC stems from nontrivial interfacial magnetic states, enabling it to persist not only under on-site disorder but also in systems with moderate lattice misalignment (see Section S5 in the Supporting Information) or an imperfect turn number in the nanoscroll (see Section S6 in the Supporting Information). It is expected that the insights and effects we have unveiled in our work will be observed in the experimental field.
Supporting Information:
Example: 1H NMR spectra for all compounds” or “Additional experimental details, materials, and methods, including photographs of experimental setup
Author Contributions
Y.-J. Z. and J.-C. Li contributed equally to this work.